OpenStudy (anonymous):

I give medals! Which is the correct simplification of 5.4*10^12/1.2*10^3 written in scientific notation? a. 4.5*10^7 b. 4.5*10^9 c. 45*10^6 d. 6.48*10^7

OpenStudy (nirmalnema):

option b is correct. 5.4/1.2 * 10^(12-3) =4.5*10^9

OpenStudy (anonymous):

\[\frac{ 5.4\times10^{12} }{ 1.2\times10^3}\]

OpenStudy (anonymous):

thanks! @nirmalnema

OpenStudy (nirmalnema):

yup.. if you take 10^3 in the numenator it will become -3

OpenStudy (nirmalnema):

your welcome..

OpenStudy (unklerhaukus):

\[5.4\times10^{12}/1.2\times10^3=\tfrac{5.4}{1.2}\times10^{12-3}\]